OpenStudy (anonymous):

What is the equation in point−slope form of the line passing through (−3, −4) and (0, 2)?

OpenStudy (michele_laino):

the requested slope is given by the subsequent formula: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 4 - 2}}{{ - 3 - 0}} = ...?\]